using function: f(x) = 3x -5
what does f(x+4)=

Answer by checkley77(12844)

f(x) = 3x -5
what does f(x+4)=
f(x+4)=3(x+4)-5
f(x+4)=3x+12-5
f(x+4)=3x+7